OpenStudy (anonymous):

The roots of a quadratic equation are 4 and -5. Which quadratic equation has these roots?

OpenStudy (anonymous):

1) (x-4) (x+5)=0 2) (x+4) (x-5)=0 3)(x-4) (x-5)=0 4) (x+4) (x+5)=0

hartnn (hartnn):

if 'a' is a root of a quadratic equation, then \(\large (x-a)\) is one of the factor! so if '4' is a root, what is the factor ?